OpenStudy (anonymous):

riza515 Group Title The U.S. quarter has a mass of 5.67 g and is approximately 1.55 mm thick.How many quarters would have to be stacked to reach 575 ft, the height of the Washington Monument?How much would this stack weigh?How much money would this stack contain?How much money would this stack contain?The US National Debt Clock showed the outstanding public debt to be $11,687,233,914,811.11 on August 19, 2009. How many stacks like the one described would be necessary to pay off this debt?

OpenStudy (jfraser):

the thickness of the quarter gives you a RATIO you can use as a conversion factor\[575\cancel{ft} * \frac{\cancel{inch}}{\cancel{ft}} * \frac{\cancel{mm}}{\cancel{inch}}* \frac{quarter}{\cancel{mm}}\]you just have to plug in the proper numbers. Units are more important than numbers are, so if you can start to look at the units of a problem instead of the numbers, you'll get a lot better a lot faster